Light of wavelength 725 nm in vacuum is incident o
Answer #1

As we know that

Fraunhofer Single Slit

:e 8 Condition for minimum mAD

The water changes the wavelength of the light to 725 . 10^–9 / 1.33 = 545.112782 nm

y (for third dark fringe)/D = theta = m*lambda/d

y (for third dark fringe)/D = theta =

1. n=3,

(n-0.5)λ= d*Sinθ

Sinθ= (n-0.5)*λ/d = (3-0.5)545.112782 * 10^-9 /1.6 * 10^-5 = 0.085173872

θ= 4.886023271 degrees